OpenStudy (anonymous):

Solve: 3m - 5 = 7m - 21 A. m = -2/13 B. m = 2/13 C. m = 4 D. m = 13/2

OpenStudy (theopenstudyowl):

Do u know how to solve equations?

OpenStudy (theopenstudyowl):

If not, u should learn! I can help

OpenStudy (anonymous):

um no I don't but that would be amazing

OpenStudy (anonymous):

combine like terms so the 3m and the 7m should be on the same side and the -5 and -21 should be on the other side

OpenStudy (anonymous):

and then solve for m

OpenStudy (theopenstudyowl):

3m-5=7m-21 3m-7m=-21+5 -4m=-16, m=4

OpenStudy (anonymous):

oh dear that just confused me

OpenStudy (anonymous):

sweety the answer is d